A company produces three products, x, y, and z. Each item of product x requires 20 units of steel, 2 units of plastic, and 1 unit of glass. Each item of product y requires 25 units of steel, 5 units of plastic, and no units of glass. Each item of product z requires 150 units of steel, 10 units of plastic, and 0.5 units of glass. The available amounts of steel, plastic and glass are, respectively, 2550, 290, and 34. How many items of each type can the company produce and utilize all the available raw materials?
Answer by ewatrrr(24785)


Hi
Set Up and solve the following system of equations:
20x + 25y + 150z = 2550
2x + 5y + 10z = 290
x + .5z = 34
x = 30
y = 30
z = 8
CHECKING our Answer***
600 + 750 + 1200 = 2550
60 + 150 + 80 = 290
30 + 4 = 34